OpenStudy (anonymous):

Simplify and show your work: |3(-2) - 7| - 7 = ?

OpenStudy (anonymous):

Using the order of operations. First solve within the absolute value. |3(-2)-7 |=| -6-7 | = |-13| = 13 13-7 = 6